# Subvarieties of pointed Abelian $\ell$-groups

### v1

@@ 6 Applying Mundici functor
${{\mathbb{K}}}=\mathbf{ISPP}_{\mathrm{U}}(\{{\boldsymbol{\mathit{Z}}}_{i}\mid i\in I\}\cup\{{\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{d_{j}}\mid j\in J,d_{j}\in D_{j}\}\cup\{\mathbf{S}_{k}\mid k\in K\})$
In the rest of the section we will focus on proving Lemma 6.5. Before doing so, we need a few smaller lemmas.
Lemma 6.10.
Let ${\boldsymbol{\mathit{A}}}_{u}$ be a totally ordered strongly positively pointed Abelian $\ell$-group. Then for each $a\in A$ there is a unique $n\in\mathbb{Z}$ and unique $r\in A$ such that $0\leq r<u$ and $a=n\cdot u+r$.
Proof.
Since $u$ is strong unit of $\boldsymbol{\mathit{A}}$} there exist $n,m\in\mathbb{Z}$ such that $n\cdot u\leq a<m\cdot u$. Let us fix maximum such $n$. Then we have $n\cdot u\leq a<(n+1)\cdot u$ and $0\leq a-n\cdot u<u$. Let us set $r=a-n\cdot u$. We indeed have $n\cdot u+r=a$.
We prove uniqueness of such decomposition. Assume $a=n_{1}\cdot u+r_{1}=n_{2}\cdot u+r_{2}$ for some $n_{1},n_{2}\in\mathbb{Z}$ and $r_{1},r_{2}\in A$ such that $0\leq r_{1},r_{2}<u$. We get $(n_{1}-n_{2})\cdot u=r_{2}-r_{1}$. Since $u<r_{2}-r_{1}<u$ we obtain $n_{1}=n_{2}$ and $r_{1}=r_{2}$. This completes the proof. ∎
The following observation directly follows from the definition of $\oplus$ and $\otimes$.
Lemma 6.11.
Let ${\boldsymbol{\mathit{A}}}_{u}$ be a positively pointed Abelian $\ell$-group and $0\leq a,b\leq u$.
  1. 1.
    If $a+b\leq u$ or $a\oplus b<u$ then $a+b=a\oplus b$.
  2. 2.
    If $a+b\geq u$ or $a\otimes b>0$ then $a+b-u=a\otimes b$.
Lemma 6.12.
Let $\boldsymbol{\mathit{A}}$} be a totally ordered pointed Abelian $\ell$-group and let $F$ be a finite subset of $A$. Let $f_{F}$ be a partial embedding of $F$ into a class of pointed Abelian $\ell$-groups $\mathbb{K}$}}. Let $a,b\in F$ such that $a\leq b$. Then $f_{F}(a)\leq f_{F}(b)$. Moreover, if $a\neq b$, then $f_{F}(a)<f_{F}(b)$.
Proof.
Since $a\lor b=b$ and $a,b,a\lor b\in F$ we have by definition of partial embedding that $f_{F}(a)\lor f_{F}(b)=f_{F}(a\lor b)$, thus $f_{F}(a)\leq f_{F}(a\lor b)=f_{F}(b)$. Also, whenever $a\neq b$ then, since $f_{F}$ is injective, we obtain $f_{F}(a)<f_{F}(b)$. ∎
Now, we can finally prove the Lemma 6.5.
Proof of Lemma 6.5.
Using Lemma 2.3 we have to show that ${\boldsymbol{\mathit{A}}}_{u}$ is partially embeddable into $\mathbb{K}$}} iff $\Gamma({\boldsymbol{\mathit{A}}}_{u})$ is partially embeddable into $\Gamma[{{\mathbb{K}}}]$.
First assume that for strongly pointed $\ell$-group ${\boldsymbol{\mathit{A}}}_{u}$ there is a family of partial embeddings $\{f_{F}\}_{F\subseteq A,|F|<\omega}$, where $f_{F}$ is a partial embedding of the set $F$ into some ${\boldsymbol{\mathit{B}}}_{v}\in{{\mathbb{K}}}$.
Let $G\subseteq\Gamma({\boldsymbol{\mathit{A}}}_{u})$ be an arbitrary finite set. Let us define $H=\{g_{1}+g_{2},-g_{1}\mid g_{1},g_{2}\in G\}$. We show that the mapping $f_{H}$ is a partial embedding of $G$ into $\Gamma({\boldsymbol{\mathit{B}}}_{v})\in\Gamma[{{\mathbb{K}}}]$. We already know that $f_{H}$ is injective and it preserves constants. Since $\otimes,\lor$ and $\land$ are definable in MV-algebras using the other connectives it is enough to show that $f_{H}$ preserves $\oplus$ and $\neg$.
We show $f_{H}$ preserves $\oplus$. Let us assume $a_{1},a_{2},a_{1}\oplus a_{2}\in G$. Let us consider two cases:
  1. 1.
    Assume $a_{1}+a_{2}\geq u$. Then we have by Lemma 6.12 $f_{H}(a_{1}+a_{2})\geq f_{H}(u)=v$ and thus
    $f_{H}(a_{1}\oplus a_{2})=f_{H}((a_{1}+a_{2})\land u)=f_{H}(u)=v=\\ f_{H}(a_{1}+a_{2})\land v=f_{H}(a_{1})+f_{H}(a_{2})\land v=f_{H}(a_{1})\oplus f_{H}(a_{2}).$
  2. 2.
    Now assume $a_{1}+a_{2}\leq u$. Then we have by Lemma 6.12 $f_{H}(a_{1}+a_{2})\leq f_{H}(u)=v$
    We have
    $f_{H}(a_{1}\oplus a_{2})=f_{H}((a_{1}+a_{2})\land u)=f_{H}(a_{1}+a_{2})=\\ f_{H}(a_{1})+f_{H}(a_{2})=(f_{H}(a_{1})+f_{H}(a_{2}))\land v=f_{H}(a_{1})\oplus f_{H}(a_{2}).$
It remains to check $f_{H}$ preserves $\neg$. Let us assume $a,\neg a\in G$. Then $a\in H$ and we have
$f_{H}(\neg a)=f_{H}(u-a)=f_{H}(u)+f_{H}(-a)=v-f_{H}(a)=\neg f_{H}(a).$
Therefore, $f_{H}$ preserves $\neg$. This shows $f_{H}$ is indeed a partial embedding of $G$ from $\Gamma({\boldsymbol{\mathit{A}}})$ into $\Gamma({\boldsymbol{\mathit{B}}})$. Since $G$ was arbitrary set we obtain that $\Gamma({\boldsymbol{\mathit{A}}})$ is partially embeddable into $\Gamma({\boldsymbol{\mathit{B}}})$.
Now, we will prove that if $\Gamma({\boldsymbol{\mathit{A}}}_{u})$ is partially embeddable into $\Gamma({\boldsymbol{\mathit{B}}}_{v})$ then also ${\boldsymbol{\mathit{A}}}_{u}$ is partially embeddable into ${\boldsymbol{\mathit{B}}}_{v}$.
Assume that $\Gamma({\boldsymbol{\mathit{A}}}_{u})$ is partially embeddable into $\Gamma({\boldsymbol{\mathit{B}}}_{v})$. Therefore for any finite set $S\subseteq\Gamma({\boldsymbol{\mathit{A}}}_{u})$ there is partial embedding $g_{S}:S\rightarrow\Gamma({\boldsymbol{\mathit{B}}}_{v})$.
Let $F\subseteq A_{u}$ be finite. Let us define a finite superset $F^{\prime}$, which is closed under - and also that $(n\cdot u+a)\in F^{\prime}\Rightarrow(n\cdot u)\in F^{\prime}\land a\in F^{\prime}$ for $n\in{\boldsymbol{\mathit{Z}}}$ and $a\in[0,u)$. We denote $F_{0}=F^{\prime}\cap[0,u]$. For any $a\in A$ we denote the decomposition from Lemma 6.10 $a=n_{a}\cdot u+r_{a}$, where $n_{a}\in Z$ and $r_{a}\in[0,u)$. Since $F_{0}\subseteq\Gamma({\boldsymbol{\mathit{A}}}_{u})$ we can define a partial a embedding $f_{F}\rightarrow B_{v}$ as follows:
$f_{F}(a)=(n_{a}\cdot v+g_{F_{0}}(r_{a})).$
This mapping is clearly injective, since we can construct $a$ knowing $n_{a},r_{a}$. We will verify it is indeed a partial homomorphism. Let $a,b,a+b\in F$. We have
$f_{F}(a+b)=n_{a+b}\cdot v+g_{F_{0}}(r_{a+b}).$
Now let us distinguish two cases.
  1. 1.
    First let us consider the case $r_{a}+r_{b}<u$ and thus by Lemma 6.11 we obtain $r_{a+b}=r_{a}+r_{b}=r_{a}\oplus r_{b}$ and $n_{a+b}=n_{a}+n_{b}$. By Lemma 6.12 and 6.11 we have
    $v>g_{F_{0}}(r_{a}\oplus r_{b})=g_{F_{0}}(r_{a})\oplus g_{F_{0}}(r_{b})=g_{F_{0}}(r_{a})+g_{F_{0}}(r_{b})$
    and thus
    $f_{F}(a+b)=n_{a}\cdot v+n_{b}\cdot v+g_{F_{0}}(r_{a}+r_{b})=n_{a}\cdot v+n_{b}\cdot v+g_{F_{0}}(r_{a}\oplus r_{b})=$
    $=n_{a}\cdot v+n_{b}\cdot v+g_{F_{0}}(r_{a})+g_{F_{0}}(r_{b})=$
    $n_{a}\cdot v+g_{F_{0}}(r_{a})+n_{b}\cdot v+g_{F_{0}}(r_{b})=f_{F}(a)+f_{F}(b).$
  2. 2.
    Now consider the case when $r_{a}+r_{b}>u$. Then we have $r_{a}+r_{b}=u+r_{a+b}$, $n_{a}+n_{b}+1=n_{a+b}$ and by Lemma 6.11 $r_{a}\otimes r_{b}=r_{a}+r_{b}-u$. By Lemma 6.12 and 6.11 we get
    $0<g_{F_{0}}(r_{a}\otimes r_{b})=g_{F_{0}}(r_{a})\otimes g_{F_{0}}(r_{b})=g_{F_{0}}(r_{a})+g_{F_{0}}(r_{b})-v$
    and thus
    $f_{F}(a+b)=(n_{a}+n_{b}+1)\cdot v+g_{F_{0}}(r_{a}+r_{b}-u)=$
    $(n_{a}+n_{b}+1)\cdot v+g_{F_{0}}(r_{a}\otimes r_{b})=$
    $(n_{a}+n_{b})\cdot v+v+g_{F_{0}}(r_{a})+g_{F_{0}}(r_{b})-v=$
    $n_{a}\cdot v+g_{F_{0}}(r_{a})+n_{b}\cdot v+g_{F_{0}}(r_{b})=f_{F}(a)+f_{F}(b).$
  3. 3.
    In the last case $r_{a}+r_{b}=u$ we have $r_{a+b}=0$, $n_{a+b}=n_{a}+n_{b}+1$ and $g_{F_{0}}(r_{a}+r_{b})=v$. Therefore,
    $f_{F}(a+b)=(n_{a}+n_{b}+1)\cdot v=$
    $n_{a}\cdot v+n_{b}\cdot v+g_{F_{0}}(r_{a}+r_{b})=n_{a}\cdot v+n_{b}\cdot v+g_{F_{0}}(r_{a}\oplus r_{b})=$
    $n_{a}\cdot v+g_{F_{0}}(r_{a})+n_{b}\cdot v+g_{F_{0}}(r_{b})=f_{F}(a)+f_{F}(b).$
This shows $f_{F}$ is preserving ++. Clearly, $f_{F}$ is preserving 0 and $a$. We will show $f_{F}$ is preserving -.
We have $f_{F}(a)=-(n_{a}\cdot v+g_{F_{0}}(r_{a}))=-n_{a}\cdot v-g_{F_{0}}(r_{a})$. If $r_{a}=0$ we have $f_{F}(a)=-n_{a}\cdot v=n_{-a}\cdot v=f_{F}(-a)$. Otherwise we have $n_{a}\cdot v-g_{F_{0}}(r_{a})<-n_{a}\cdot v$ and thus $n_{a}=n_{-a}+1$ and $u-r_{a}=r_{-a}$. By applying $g_{F_{0}}$ we get
$v-g_{F_{0}}(r_{a})=\neg g_{F_{0}}(r_{a})=g_{F_{0}}(\neg r_{a})=g_{F_{0}}(u-r_{a})=g_{F_{0}}(r_{-a}).$
Therefore, we have
$f_{F}(a)=-n_{a}\cdot v-g_{F_{0}}(r_{a})=n_{-a}\cdot v+v-v+g_{F_{0}}(r_{-a})=n_{-a}\cdot v+g_{F_{0}}(r_{-a})=f_{F}(-a).$
This shows $f_{F}$ is preserving -. It remains to check $f_{F}$ preserves lattice operations. Since $\boldsymbol{\mathit{A}}$} is totally ordered it is enough to check $f_{F}$ preserves ordering. Let $a\leq b$ and consequently $n_{a}\leq n_{b}$. We distinguish two cases
  1. 1.
    If $n_{a}=n_{b}$ then $r_{a}\leq r_{b}$. By Lemma 6.12, also $g_{F_{0}}(r_{a})\leq g_{F_{0}}(r_{b})$ and therefore we have the following:
    $f_{F}(a)=n_{a}\cdot v+g_{F_{0}}(r_{a})\leq n_{a}\cdot v+g_{F_{0}}(r_{b})=f_{F}(b).$
  2. 2.
    If $n_{a}<n_{b}$, by using Lemma 6.12 we obtain $g_{F_{0}}(r_{a})\leq g_{F_{0}}(u)=v$ and $0=g_{F_{0}}(0)\leq g_{F_{0}}(r_{b})$ and thus we have
    $f_{F}(a)=n_{a}\cdot v+g_{F_{0}}(r_{a})\leq(n_{a}+1)\cdot v\leq n_{b}\cdot v\leq n_{b}\cdot v+g_{F_{0}}(r_{b})=f_{F}(b).$
This shows $f_{F}$ preserves ordering and thus $f_{F}$ a partial embedding. Since the set $F$ was arbitrary, we conclude that ${\boldsymbol{\mathit{A}}}_{u}$ is partially embeddable into $\mathbb{K}$}}. ∎
$p\mathbb{AL}$$\mathcal{V}(\mathbf{Z}_{0})$$\mathcal{T}$$p\mathbb{AL}^{+}$$\mathcal{V}(\mathbf{Z}_{2})$$\mathcal{V}(\mathbf{Z}_{3})$$\mathcal{V}(\mathbf{Z}_{5})$$\mathcal{V}(\mathbf{Z}_{4})$$\mathcal{V}(\mathbf{Z}_{6})$$\mathcal{V}(\mathbf{Z}_{9})$$\mathcal{V}(\mathbf{Z}_{12})$$\mathcal{V}(\mathbf{Z}_{2}\overrightarrow{\times}\mathbf{Z}_{0})$$\mathcal{V}(\mathbf{Z}_{3}\overrightarrow{\times}\mathbf{Z}_{0})$$\mathcal{V}(\mathbf{Z}_{5}\overrightarrow{\times}\mathbf{Z}_{0})$$\mathcal{V}(\mathbf{Z}_{4}\overrightarrow{\times}\mathbf{Z}_{0})$$\mathcal{V}(\mathbf{Z}_{6}\overrightarrow{\times}\mathbf{Z}_{0})$$\mathcal{V}(\mathbf{Z}_{12}\overrightarrow{\times}\mathbf{Z}_{0})$$\mathcal{V}(\mathbf{Z}_{9}\overrightarrow{\times}\mathbf{Z}_{0})$$p\mathbb{AL}^{-}$$\mathcal{V}(\mathbf{Z}_{-2})$$\mathcal{V}(\mathbf{Z}_{-3})$$\mathcal{V}(\mathbf{Z}_{-5})$$\mathcal{V}(\mathbf{Z}_{-4})$$\mathcal{V}(\mathbf{Z}_{-6})$$\mathcal{V}(\mathbf{Z}_{-9})$$\mathcal{V}(\mathbf{Z}_{-12})$$\mathcal{V}(\mathbf{Z}_{-2}\overrightarrow{\times}\mathbf{Z}_{0})$$\mathcal{V}(\mathbf{Z}_{-3}\overrightarrow{\times}\mathbf{Z}_{0})$$\mathcal{V}(\mathbf{Z}_{-5}\overrightarrow{\times}\mathbf{Z}_{0})$$\mathcal{V}(\mathbf{Z}_{-4}\overrightarrow{\times}\mathbf{Z}_{0})$$\mathcal{V}(\mathbf{Z}_{-6}\overrightarrow{\times}\mathbf{Z}_{0})$$\mathcal{V}(\mathbf{Z}_{-12}\overrightarrow{\times}\mathbf{Z}_{0})$$\mathcal{V}(\mathbf{Z}_{-9}\overrightarrow{\times}\mathbf{Z}_{0})$
Figure 1: The lattice of subvarieties of $p\mathbb{AL}$.
[…]
Figure 2: The lattice of subvarieties of $p\mathbb{AL}$

### v2

@@ 6.1 Gispert’s classification of universal classes of totally ordered MV-algebras
Let $\mathbb{K}$}} be the quasivariety generated by a class of totally ordered pointed Abelian $\ell$-groups. Then there exists $I,J,K\subseteq\mathbb{Z}$ and for every $j\in J\setminus\{0\}$, a nonempty subset $D_{j}\subseteq\operatorname{div}(j)$ such that
7 Future Work
In this section, we outline possible directions for future research on the topics covered in this paper.
  • In this paper, we discuss the characterization of all subquasivarieties generated by totally ordered $\ell$-groups. However, we provide only the structural characterization of these subquasivarieties without any quasiequational basis. The obvious future step is to provide these axiomatizations, which we can obtain by analyzing the algebras ${\boldsymbol{\mathit{Z}}}_{i}$, ${\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{d_{j}}$ and $\mathbf{S}_{k}$ from Corollary 6.9.
  • One of the main motivations for this paper was to show that when we discuss the subvarieties of $p\mathbb{AL}$, we can do so purely in the theory of pointed Abelian $\ell$-groups in a self-contained way without any circular reasoning. However, the same still needs to be done regarding the characterization of subquasivarieties of $p\mathbb{AL}$ that are generated by their totally ordered members.
  • One of the most significant drawbacks of Corollary 3.12 is that we require the strongly pointed $\ell$-group to be totally ordered. It is known that there are subquasivarieties of $p\mathbb{AL}$ that are not generated by their totally ordered members and they have no immediate counterpart in MV-algebras. For example, consider the quasivariety given by the quasiequation $\mathtt{f}\geq 0\implies x\geq 0$ (see [9]), which is known to not be generated by its totally ordered members. We do not aim to provide an explicit description of all subquasivarieties of $p\mathbb{AL}$, as such a class is likely Q-universal, similarly to the class of all subquasivarieties of MV-algebras [1]. Instead, we aim to explicitly describe the additional subquasivarieties that do not correspond to any subquasivariety of MV-algebras.
  • An additional question is connected to expanding the language of pointed Abelian $\ell$-groups with a modal operator $\langle{{\boldsymbol{\mathit{A}}}_{a},\Box}\rangle$ [22]. By doing so, we obtain a variety of modal pointed Abelian $\ell$-groups that is not generated by totally ordered algebras. As such, we cannot utilize Theorem 3.10 here, meaning we need to reformulate more general criteria for when a pointed modal Abelian $\ell$-group is in a variety generated by its strongly pointed modal convex $\ell$-subgroup.